OpenStudy (anonymous):

A country's population in 1991 was 147 million. In 1998 it was 153 million. Estimate the population in 2017 using the exponential growth formula. Round your answer to the nearest million. P = Aekt

OpenStudy (anonymous):

e^kt = P/A kt = ln(P/A) k = ln(P/A) / t plugging in, k = ln(153/147) / 7 = 0.0057 to 4 dp P(26) = 147e^(0.0057*26) = 170 mn <-------

OpenStudy (anonymous):

Is It Correct

OpenStudy (anonymous):

Thank you @Pizza7 Yes it is!

OpenStudy (anonymous):

k Give Me The Best Response Button Please

OpenStudy (anonymous):

Yeahhh Thank YOU!!!